OpenStudy (anonymous):

Please help me solve 5 1/6 - 2 2/3=

OpenStudy (anonymous):

31/6 - 8/3= 31/6 - 16/3= 15/3= 5

OpenStudy (anonymous):

omg, yo're right, i'm sorry

OpenStudy (anonymous):

its fyn. you got the method dude. that is what matters.

OpenStudy (anonymous):

i'm sorry

OpenStudy (anonymous):

5 1/6 - 2 2/3 31/6 - 8/3 31/6 - 16/6 (LCD) = 15/6 or 2 1/2

OpenStudy (anonymous):

@agit you method is right but calculation mistake. 31/6 - 8/3= 31/6 - 16/6= 15/6 is tha answer. 2 1/2 is tha answer.